A gardener wants to create a triangular garden in the shape of a right triangle with shortest side length x – 4y ft. and the middle length side x + 6y ft. What is an algebraic expression for the area of the garden?
Answer by nerdybill(6948)

A gardener wants to create a triangular garden in the shape of a right triangle with shortest side length x – 4y ft. and the middle length side x + 6y ft. What is an algebraic expression for the area of the garden?
.
area = (1/2)base*height
area = (1/2)(x-4y)(x+6y)
area = (1/2)(x^2+6xy-4xy-24y^2)
area = (1/2)(x^2+2xy-24y^2)
area = (x^2)/2+xy-12y^2